# Turbocharging Gaussian Process Inference with Approximate Sketch-and-Project

### v1

@@ 4.2 ADASAP algorithm
     # Phase I: Compute matrix-matrix product $K_{\mathcal{B}n}W_{t}$triangleright Costs $\mathcal{O}(bnn_{\textrm{rhs}}/\mathrm{N_{workers}})$
@@
     $K_{\mathcal{B}n}W_{t}\leftarrow\texttt{ColDistMatMat}(W_{t},\mathcal{B},\{\mathcal{W}_{1},\ldots,\mathcal{W}_{\mathrm{N_{workers}}}\})$triangleright Algorithm 3
     # Phase II: Compute Nyström sketch $K_{\mathcal{B}\mathcal{B}}\Omega$triangleright Costs $\mathcal{O}(rb^{2}/\mathrm{N_{workers}}$)
@@
     $K_{\mathcal{B}\mathcal{B}}\Omega\leftarrow\texttt{RowDistMatMat}(\Omega,\mathcal{B},\{\mathcal{W}_{1},\ldots,\mathcal{W}_{\mathrm{N_{workers}}}\})$triangleright Algorithm 4
     # Phase III: Compute Nyström approximation and get stepsize $\triangleright$ Costs $\mathcal{O}(b^{2}+br^{2})$
     $U,S\leftarrow\texttt{RandNysAppx}(K_{\mathcal{B}\mathcal{B}}\Omega,\Omega,r)$triangleright Algorithm 5
     $P_{\mathcal{B}}\leftarrow USU^{T}+(S_{rr}+\lambda)I$triangleright Get preconditioner. Never explicitly formed!
     $\eta_{\mathcal{B}}\leftarrow\texttt{GetStepsize}(P_{\mathcal{B}},K_{\mathcal{B}\mathcal{B}}+\lambda I)$triangleright Algorithm 6
     # Phase IV: Compute updates using acceleration $\triangleright$ Costs $\mathcal{O}(brn_{\textrm{rhs}}+nn_{\textrm{rhs}})$
@@
     $W_{t+1},V_{t+1},Z_{t+1}\leftarrow\texttt{NestAcc}(W_{t},V_{t},Z_{t},D_{t},\eta_{\mathcal{B}},\beta,\gamma,\alpha)$triangleright Algorithm 7
@@
return $\bar{W}_{T}$ if tail_average else $W_{T}$triangleright Approximate solution to $K_{\lambda}W=Y$
The pseudocode for ADASAP is presented in Algorithm 2. ADASAP uses tail averaging, approximate subspace preconditioning, distributed matrix-matrix products, and Nesterov acceleration. Assuming perfect parallelism, ADASAP has per iteration runtime of $\mathcal{O}(nn_{\textrm{rhs}}b/\mathrm{N_{workers}}+b^{2}r/\mathrm{N_{workers}})$, a significant improvment over the $\mathcal{O}(nn_{\textrm{rhs}}b+b^{3})$ iteration time of SAP. The per-iteration runtime of ADASAP is comparable to distributed SDD—in other words, ADASAP effectively preconditions the problem, reducing the total iterations required, without significantly slowing down each iteration. Moreover, unlike SDD, ADASAP automatically sets the stepsize at each iteration, removing the need for expensive tuning. As a default, we set the acceleration parameters $\mu$ and $\nu$ to $\lambda$ and $n/b$, respectively, as in Rathore et al. (2025), and find they yield excellent performance in Section 5.
@@
Our theory in Section 3 does not cover the approximate preconditioning, acceleration, or uniform sampling that is used in ADASAP. Despite the theory-practice gap between SAP (Algorithm 1) and ADASAP (Algorithm 2), we believe the theory developed for SAP can be extended to ADASAP. This belief is rooted in Dereziński and Yang (2024); Dereziński et al. (2024); Rathore et al. (2025), which show that SAP methods still converge when using approximate preconditioning, acceleration, and uniform sampling. We leave this extension as a direction for future research.
For simplicity, our experiments run ADASAP without tail averaging. Tail averaging is needed to establish Theorem 3.2, but we do not expect it to yield significant practical improvements. This is in line with the SGD literature, where averaging leads to better theoretical convergence rates, but the last iterate delivers similar performance in practice (Shamir and Zhang, 2013; Johnson and Zhang, 2013).
@@ 5 Experiments
We present three sets of experiments showing ADASAP outperforms state-of-the-art methods for GP inference: (i) GP inference on large benchmark datasets (Section 5.1), (ii) GP inference on huge-scale transportation data analysis with $n>3\cdot 10^{8}$ samples (Section 5.2), and (iii) the Bayesian optimization task from Lin et al. (2023) (Section 5.3). We evaluate ADASAP against the following competitors:
@@
Our experiments are run in single precision on 48 GB NVIDIA RTX A6000 GPUs using Python 3.10, PyTorch 2.6.0 (Paszke et al., 2019), and CUDA 12.5. We use 2, 3, and 1 GPU(s) per experiment in Sections 5.1, 5.2 and 5.3, respectively. Code for reproducing our experiments is available at https://github.com/pratikrathore8/scalable_gp_inference. Additional details are in Appendix D.
@@ 5.1 GP inference on large-scale datasets
We benchmark on six large-scale regression datasets from the UCI repository, OpenML, and sGDML (Chmiela et al., 2017). The results are reported in Table 1. ADASAP achieves the lowest RMSE and mean negative log-likelihood (NLL) on each dataset; the NLL is computed using 64 posterior samples (via pathwise conditioning). SDD-10 is competitive with ADASAP on yolanda and acsincome, but performs much worse on the other datasets. SDD is also sensitive to the stepsize: SDD-1 attains a larger RMSE and NLL than ADASAP on all datasets, while SDD-100 diverges on all datasets. PCG obtains the same RMSE and NLL as ADASAP on yolanda and song, but its performance degrades for larger datasets. Additionally, Fig. 2 shows ADASAP achieves the lowest RMSE and NLL throughout the optimization process, demonstrating its efficiency with respect to wall-clock time.
@@
5.2 Showcase: Transportation data analysis with $>3\cdot 10^{8}$ samples
To demonstrate the power of ADASAP on huge-scale problems, we perform GP inference on a subset of the taxi dataset (https://github.com/toddwschneider/nyc-taxi-data) with $n=3.31\cdot 10^{8}$ samples and dimension $d=9$: the task is to predict taxi ride durations in New York City. To the best of our knowledge, this is the first time that full GP inference has been scaled to a dataset of this size. Due to memory constraints, we are unable to compute 64 posterior samples as done in Section 5.1 for computing NLL, so we only report RMSE.
The results are shown in Fig. 3. ADASAP obtains the lowest RMSE out of all the methods. Once again, SDD is sensitive to the stepsize: SDD-1 obtains an RMSE of 0.60, as opposed to ADASAP, which obtains an RMSE of 0.50, while SDD-100 diverges. SDD-10 obtains an RMSE of 0.52, which is similar to that of ADASAP. However, ADASAP reaches the RMSE attained by SDD-10 in 45% less time than SDD-10, which translates to a difference of 14 hours of runtime! Furthermore, PCG runs out of memory, as the memory required for storing the sketch in single precision is $3.31\cdot 10^{8}\cdot 100\cdot 4\approx 130$ GB, which exceeds the 48 GB of memory in the A6000 GPUs used in our experiments.
Figure 3: Comparison between ADASAP and competitors on transportation data analysis. ADASAP attains the lowest RMSE and it obtains a 1.8×1.8$\times$1.8 × speed up over the second-best method, SDD-10. SDD-100 diverges and PCG runs out of memory, so they do not appear in the figure.
@@ 5.3 Large-scale Bayesian optimization
The results are shown in Table 2. ADASAP makes the biggest improvement over the random search baseline. As we have seen in the other experiments, SDD is sensitive to the stepsize: SDD-10 and SDD-100 make no progress. PCG also makes less progress than ADASAP across both lengthscales.
@@ 6 Conclusion
We introduce ADASAP, an approximate, distributed, accelerated sketch-and-project method for GP inference. We demonstrate that ADASAP outperforms state-of-the-art GP inference methods like PCG and SDD on large-scale benchmark datasets, a huge-scale dataset with $>3\cdot 10^{8}$ samples, and large-scale Bayesian optimization. Moreover, we show that SAP-style methods are theoretically principled for GP inference—we prove that SAP is the first efficient, condition number-free algorithm for estimating the posterior mean along the top spectral basis functions. Future work should extend the theoretical results for SAP to ADASAP and investigate ADASAP in lower precision (e.g., float16).
@@ A.2 Sketch-and-project
In addition to improvements in the analysis of exact sketch-and-project for linear systems, significant strides have been made when approximate subspace preconditioning is employed. Dereziński and Yang [2024], Dereziński et al. [2024] both use PCG to approximately solve the linear system defined by the subspace Hessian in each iteration of SAP, while Dereziński et al. [2025] augment that system with Tikhonov regularization to further improve its conditioning. Rathore et al. [2025] consider an approximate sketch-and-project method for solving kernel ridge regression with one right-hand side. They replace $K_{\mathcal{B}\mathcal{B}}+\lambda I$ in SAP with $\hat{K}_{\mathcal{B}\mathcal{B}}+\rho I$, where $\hat{K}_{\mathcal{B}\mathcal{B}}$ is a randomized Nyström approximation of $K_{\mathcal{B}\mathcal{B}}$. They establish linear convergence for the method, with a rate that is comparable to exact SAP when the kernel matrix exhibits an appropriate rate of spectral decay. Of the approximate SAP methods, ADASAP is closest to Rathore et al. [2025], but with significant differences such as tail averaging, handling multiple right-hand sides, and distributing bottleneck operations over multiple devices.
To the best of our knowledge, our work is the first to systematically investigate the convergence rate of SAP with tail averaging along the top eigenspace. Epperly et al. [2024a] combines tail averaging with randomized Kaczmarz for solving inconsistent linear systems, but they do not investigate the convergence rate along the top spectral subspaces. Steinerberger [2021], Dereziński and Rebrova [2024] look at the convergence rate of the expected iterates along a particular eigenvector. However, as noted in Section 3.2, this does not lead to a meaningful convergence rate for the iterates produced by the algorithm, and this also fails to capture the behavior over an entire subspace. Thus, our analysis provides the first concrete characterization of the convergence rate of tail-averaged SAP along the top eigenvectors.
@@ B.1 Overview of the analysis
Let $w_{\star}:=K_{\lambda}^{-1}y\in\mathbb{R}^{n}$ be the solution of the linear system defined by $K_{\lambda}$ and $y$, an let us use $w_{t}\in\mathbb{R}^{n}$ to denote the iterates produced by sketch-and-project when solving that system. To convert from the posterior mean error to the Euclidean norm, we first write the spectral basis functions as $u^{(i)}=\frac{1}{\sqrt{\lambda_{i}}}k(\cdot,X)v_{i}$, where $(\lambda_{i},v_{i})$ is the $i^{th}$-largest eigenpair of $K$. Thus, for any $w\in\mathbb{R}^{n}$ and $m=k(\cdot,X)w\in\mathcal{H}$, we have $\|\mathrm{proj}_{\ell}(m-m_{n})\|_{\mathcal{H}}=\|Q_{\ell}K^{1/2}(w-w_{\star})\|\leq\|Q_{\ell}K_{\lambda}^{1/2}(w-w_{\star})\|$, where $Q_{\ell}=\sum_{i=1}^{\ell}v_{i}v_{i}^{T}$ is the projection onto the $-\ell$ subspace of $K$.
@@
where we used that $\bar{\Pi}$, $K_{\lambda}$, and $Q_{\ell}$ commute with each other. This would suggest that we should obtain a fast linear rate $(1-\frac{1}{1+\phi(b,\ell)})^{t}$ for the sketch-and-project iterates in the $-\ell$ subspace norm. However, it turns out that the actual residual does not attain the same convergence guarantee as its expectation because, unlike $\bar{\Pi}$, the random projection $\Pi_{\mathcal{B}}$ does not commute with $K_{\lambda}$ or $Q_{\ell}$. Indeed,
@@
and even though the matrix $(I-\Pi_{\mathcal{B}})Q_{\ell}(I-\Pi_{\mathcal{B}})$ is low-rank, its expectation may not be, since $\Pi_{\mathcal{B}}$ does not commute with $Q_{\ell}$. This means that we have no hope of showing a linear convergence rate along the $-\ell$ subspace. Nevertheless, we are still able to obtain the following bound:
@@
This results in a bias-variance decomposition: the first term is the bias, which exhibits fast convergence in $-\ell$ subspace, while the second term is the variance, which accounts for the noise coming from the orthogonal complement subspace. We then use tail averaging to decay this noise, which allows us to benefit from the fast convergence of expected iterates through a sublinear rate $\frac{\phi(b,\ell)}{t}$.
Finally, to attain the asymptotically faster linear rate for the residual vectors, we observe that every $-\ell$ subspace norm is upper-bounded by the full norm: $\|Q_{\ell}\Delta_{t}\|\leq\|\Delta_{t}\|$. Thus, we can effectively repeat the above analysis with $\ell=n$ (since $Q_{n}=I$), in which case the variance term becomes zero, and we recover a linear rate of the form $\left(1-\frac{1}{1+\phi(b,n)}\right)^{t}$.
B.2 Convergence along $-\ell$ dimensional subspace for positive-definite matrices
In this section, we consider the general problem of solving a linear system $Aw=y$ for an $n\times n$ positive-definite matrix $A$. We provide theoretical results providing the fast convergence guarantees for the iterate sequence generated by SAP along the $-\ell$ dimensional subspace of $A$, with Theorem B.3 being the main result of this section. We use Theorem B.3 in the next section to derive the subspace convergence result for posterior mean in GP inference, proving Theorem 3.2.

### v2

@@ 4.2 ADASAP algorithm
  # Step I: Compute matrix-matrix product $K_{\mathcal{B}n}W_{t}$ $\triangleright$ Costs $\mathcal{O}(bnn_{\textrm{rhs}}/\mathrm{N_{workers}})$
@@
  $K_{\mathcal{B}n}W_{t}\leftarrow\texttt{ColDistMatMat}(W_{t},\mathcal{B},\{\mathcal{W}_{1},\ldots,\mathcal{W}_{\mathrm{N_{workers}}}\})$ $\triangleright$ Algorithm˜3
  # Step II: Compute Nyström sketch $K_{\mathcal{B}\mathcal{B}}\Omega$ $\triangleright$ Costs $\mathcal{O}(rb^{2}/\mathrm{N_{workers}}$)
@@
  $K_{\mathcal{B}\mathcal{B}}\Omega\leftarrow\texttt{RowDistMatMat}(\Omega,\mathcal{B},\{\mathcal{W}_{1},\ldots,\mathcal{W}_{\mathrm{N_{workers}}}\})$ $\triangleright$ Algorithm˜4
  # Step III: Compute Nyström approximation and get stepsize $\triangleright$ Costs $\mathcal{O}(b^{2}+br^{2})$
  $U,S\leftarrow\texttt{RandNysAppx}(K_{\mathcal{B}\mathcal{B}}\Omega,\Omega,r)$ $\triangleright$ Algorithm˜5
  $P_{\mathcal{B}}\leftarrow USU^{T}+(S_{rr}+\lambda)I$ $\triangleright$ Get preconditioner. Never explicitly formed!
  $\eta_{\mathcal{B}}\leftarrow\texttt{GetStepsize}(P_{\mathcal{B}},K_{\mathcal{B}\mathcal{B}}+\lambda I)$ $\triangleright$ Algorithm˜6
  # Step IV: Compute updates using acceleration $\triangleright$ Costs $\mathcal{O}(brn_{\textrm{rhs}}+nn_{\textrm{rhs}})$
@@
  $W_{t+1},V_{t+1},Z_{t+1}\leftarrow\texttt{NestAcc}(W_{t},V_{t},Z_{t},D_{t},\eta_{\mathcal{B}},\beta,\gamma,\alpha)$ $\triangleright$ Algorithm˜7
@@
return $\bar{W}_{T}$ if tail_average else $W_{T}$ $\triangleright$ Approximate solution to $K_{\lambda}W=Y$
The pseudocode for ADASAP is presented in Algorithm˜2. ADASAP uses tail averaging, approximate subspace preconditioning, distributed matrix-matrix products, and Nesterov acceleration. Assuming perfect parallelism, ADASAP has per iteration runtime of $\mathcal{O}(nn_{\textrm{rhs}}b/\mathrm{N_{workers}}+b^{2}r/\mathrm{N_{workers}})$, a significant improvment over the $\mathcal{O}(nn_{\textrm{rhs}}b+b^{3})$ iteration time of SAP. The per-iteration runtime of ADASAP is comparable to distributed SDD—in other words, ADASAP effectively preconditions the problem, reducing the total iterations required, without significantly slowing down each iteration. Moreover, unlike SDD, ADASAP automatically sets the stepsize at each iteration, removing the need for expensive tuning. By default, we set the acceleration parameters $\mu$ and $\nu$ to $\lambda$ and $n/b$, respectively (as done in Rathore et al. (2025)), and find they yield excellent performance in Section˜5.
@@
Our theory in Section˜3 does not cover the approximate preconditioning, acceleration, or uniform sampling that is used in ADASAP. Despite the theory-practice gap between SAP (Algorithm˜1) and ADASAP (Algorithm˜2), we believe the theory developed for SAP can be extended to ADASAP. This belief is rooted in Dereziński and Yang (2024); Dereziński et al. (2024); Rathore et al. (2025), which show that SAP methods still converge when using approximate preconditioning, acceleration, and uniform sampling. We leave this extension as a direction for future research.
For simplicity, our experiments run ADASAP without tail averaging. Tail averaging is needed to establish Theorem˜3.2, but we do not expect it to yield significant practical improvements (Section˜C.6). This is in line with the SGD literature, where averaging leads to better theoretical convergence rates, but the last iterate delivers similar performance in practice (Shamir and Zhang, 2013; Johnson and Zhang, 2013).
@@ 5 Experiments
We present three sets of experiments showing ADASAP outperforms state-of-the-art methods for GP inference: (i) GP inference on large benchmark datasets (Section˜5.1), (ii) GP inference on huge-scale transportation data analysis with $n>3\cdot 10^{8}$ samples (Section˜5.2), and (iii) the Bayesian optimization task from Lin et al. (2023) (Section˜5.3). We evaluate ADASAP against the following competitors:
@@
Our experiments are run in single precision on 48 GB NVIDIA RTX A6000 GPUs using Python 3.10, PyTorch 2.6.0 (Paszke et al., 2019), and CUDA 12.5. We use 2, 3, and 1 GPU(s) per experiment in Sections˜5.1, 5.2 and 5.3, respectively. Code for reproducing our experiments is available at https://github.com/pratikrathore8/scalable_gp_inference.
Additional details are in Appendix˜D.
@@ 5.1 GP inference on large-scale datasets
We benchmark on six large-scale regression datasets from the UCI repository, OpenML, and sGDML (Chmiela et al., 2017). The results are reported in Table˜1. ADASAP achieves the lowest RMSE and mean negative log-likelihood (NLL) on each dataset; the NLL is computed using 64 posterior samples (via pathwise conditioning). SDD-10 is competitive with ADASAP on yolanda and acsincome, but performs much worse on the other datasets. SDD is also sensitive to the stepsize: SDD-1 attains a larger RMSE and NLL than ADASAP on all datasets, while SDD-100 diverges on all datasets. PCG obtains the same RMSE and NLL as ADASAP on yolanda and song, but its performance degrades for larger datasets. Additionally, Fig.˜2 shows ADASAP achieves the lowest RMSE and NLL throughout the optimization process, demonstrating its efficiency with respect to wall-clock time.
@@
5.2 Showcase: Transportation data analysis with $3\cdot 10^{8}$ samples
To demonstrate the power of ADASAP on huge-scale problems, we perform GP inference on a subset of the taxi dataset (https://github.com/toddwschneider/nyc-taxi-data) with $n=3.31\cdot 10^{8}$ samples and dimension $d=9$: the task is to predict taxi ride durations in New York City. To the best of our knowledge, this is the first time that full GP inference has been scaled to a dataset of this size. Due to memory constraints, we are unable to compute 64 posterior samples as done in Section˜5.1 for computing NLL, so we only report RMSE.
The results are shown in Fig.˜3. ADASAP obtains the lowest RMSE out of all the methods. Once again, SDD is sensitive to the stepsize: SDD-1 obtains an RMSE of 0.60, as opposed to ADASAP, which obtains an RMSE of 0.50, while SDD-100 diverges. SDD-10 obtains an RMSE of 0.52, which is similar to that of ADASAP. However, ADASAP reaches the RMSE attained by SDD-10 in 45% less time than SDD-10, which translates to a difference of 14 hours of runtime! Furthermore, PCG runs out of memory, as the memory required for storing the sketch in single precision is $3.31\cdot 10^{8}\cdot 100\cdot 4\approx 130$ GB, which exceeds the 48 GB of memory in the A6000 GPUs used in our experiments.
Figure 3: Comparison between ADASAP and competitors on transportation data analysis. ADASAP attains the lowest RMSE and it obtains a $1.8\times$ speed up over the second-best method, SDD-10. SDD-100 diverges and PCG runs out of memory, so they do not appear in the figure.
@@ 5.3 Large-scale Bayesian optimization
The results are shown in Table˜2. ADASAP makes the biggest improvement over the random search baseline. As we have seen in the other experiments, SDD is sensitive to the stepsize: SDD-10 and SDD-100 make no progress. PCG also makes less progress than ADASAP across both lengthscales.
@@ 6 Conclusion
We introduce ADASAP, an approximate, distributed, accelerated sketch-and-project method for GP inference. We demonstrate that ADASAP outperforms state-of-the-art GP inference methods like PCG and SDD on large-scale benchmark datasets, a huge-scale dataset with $3\cdot 10^{8}$ samples, and large-scale Bayesian optimization. Moreover, we show that SAP-style methods are theoretically principled for GP inference—we prove that SAP is the first efficient, condition number-free algorithm for estimating the posterior mean along the top spectral basis functions. Future work should extend the theoretical results for SAP to ADASAP and investigate ADASAP in lower precision (e.g., float16).
@@ A.2 Sketch-and-project
In addition to improvements in the analysis of exact sketch-and-project for linear systems, significant strides have been made when approximate subspace preconditioning is employed. Dereziński and Yang [2024], Dereziński et al. [2024] both use PCG to approximately solve the linear system defined by the subspace Hessian in each iteration of SAP, while Dereziński et al. [2025] augment that system with Tikhonov regularization to further improve its conditioning. Rathore et al. [2025] consider an approximate sketch-and-project method for solving kernel ridge regression with one right-hand side. They replace $K_{\mathcal{B}\mathcal{B}}+\lambda I$ in SAP with $\hat{K}_{\mathcal{B}\mathcal{B}}+\rho I$, where $\hat{K}_{\mathcal{B}\mathcal{B}}$ is a randomized Nyström approximation of $K_{\mathcal{B}\mathcal{B}}$. They establish linear convergence for the method, with a rate that is comparable to exact SAP when the kernel matrix exhibits an appropriate rate of spectral decay.
Among the above-mentioned works, ADASAP is closest to Rathore et al. [2025], but with significant algorithmic differences such as tail averaging, handling multiple right-hand sides, and distributing bottleneck operations over multiple devices. On the theoretical side, our analysis focuses on the convergence along the top $\ell$-subspace using exact SAP with tail-averaging. In contrast, Rathore et al. [2025] focus on convergence of approximate sketch-and-project (with and without acceleration) over the entire space, and as a result, they are only able to show a fast convergence guarantee under certain strong conditions on the spectral decay (effectively limiting the condition number of the problem). Consequently, the global convergence analysis of Rathore et al. [2025] does not explain the rapid initial progress that SAP-style algorithms make on test error. We believe that our two-phase analysis consisting of (i) fast condition number-free sublinear convergence, followed by (ii) a slower global convergence rate dependent upon the subspace condition number, better captures this phenomenon. In particular, our result is the first step in developing a systematic analysis of the generalization properties of SAP-style algorithms.
To the best of our knowledge, our work is the first to systematically investigate the convergence rate of SAP with tail averaging along the top eigenspace. Epperly et al. [2024a] combines tail averaging with randomized Kaczmarz for solving inconsistent linear systems, but they do not investigate the convergence rate along the top spectral subspaces. Steinerberger [2021], Dereziński and Rebrova [2024] look at the convergence rate of the expected iterates along a particular eigenvector. However, as noted in Section˜3.2, this does not lead to a meaningful convergence rate for the iterates produced by the algorithm, and this also fails to capture the behavior over an entire subspace. Thus, our analysis provides the first concrete characterization of the convergence rate of tail-averaged SAP along the top eigenvectors.
@@ Appendix B Proofs of the Main Results
Throughout, we adopt the shorthand $A_{\lambda}\coloneqq A+\lambda I$, where $A$ is a square matrix, $\lambda\in\mathbb{R}$, and $I$ denotes the identity matrix of the same size as $A$.
@@ B.1 Overview of the analysis
Let $w_{\star}:=K_{\lambda}^{-1}y\in\mathbb{R}^{n}$ be the solution of the linear system defined by $K_{\lambda}$ and $y$, an let us use $w_{t}\in\mathbb{R}^{n}$ to denote the iterates produced by sketch-and-project when solving that system. To convert from the posterior mean error to the Euclidean norm, we first write the spectral basis functions as $u^{(i)}=\frac{1}{\sqrt{\lambda_{i}}}k(\cdot,X)v_{i}$, where $(\lambda_{i},v_{i})$ is the $i^{th}$-largest eigenpair of $K$. Thus, for any $w\in\mathbb{R}^{n}$ and $m=k(\cdot,X)w\in\mathcal{H}$, we have $\|\mathrm{proj}_{\ell}(m-m_{n})\|_{\mathcal{H}}=\|Q_{\ell}K^{1/2}(w-w_{\star})\|\leq\|Q_{\ell}K_{\lambda}^{1/2}(w-w_{\star})\|$, where $Q_{\ell}=\sum_{i=1}^{\ell}v_{i}v_{i}^{T}$ is the projection onto the top-$\ell$ subspace of $K$.
@@
where we used that $\bar{\Pi}$, $K_{\lambda}$, and $Q_{\ell}$ commute with each other. This would suggest that we should obtain a fast linear rate $(1-\frac{1}{1+\phi(b,\ell)})^{t}$ for the sketch-and-project iterates in the top-$\ell$ subspace norm. However, it turns out that the actual residual does not attain the same convergence guarantee as its expectation because, unlike $\bar{\Pi}$, the random projection $\Pi_{\mathcal{B}}$ does not commute with $K_{\lambda}$ or $Q_{\ell}$. Indeed,
@@
and even though the matrix $(I-\Pi_{\mathcal{B}})Q_{\ell}(I-\Pi_{\mathcal{B}})$ is low-rank, its expectation may not be, since $\Pi_{\mathcal{B}}$ does not commute with $Q_{\ell}$. This means that we have no hope of showing a linear convergence rate along the top-$\ell$ subspace. Nevertheless, we are still able to obtain the following bound:
@@
This results in a bias-variance decomposition: the first term is the bias, which exhibits fast convergence in top-$\ell$ subspace, while the second term is the variance, which accounts for the noise coming from the orthogonal complement subspace. We then use tail averaging to decay this noise, which allows us to benefit from the fast convergence of expected iterates through a sublinear rate $\frac{\phi(b,\ell)}{t}$.
Finally, to attain the asymptotically faster linear rate for the residual vectors, we observe that every top-$\ell$ subspace norm is upper-bounded by the full norm: $\|Q_{\ell}\Delta_{t}\|\leq\|\Delta_{t}\|$. Thus, we can effectively repeat the above analysis with $\ell=n$ (since $Q_{n}=I$), in which case the variance term becomes zero, and we recover a linear rate of the form $\left(1-\frac{1}{1+\phi(b,n)}\right)^{t}$.
B.2 Convergence along top-$\ell$ dimensional subspace for positive-definite matrices
In this section, we consider the general problem of solving a linear system $Aw=y$ for an $n\times n$ positive-definite matrix $A$. We provide theoretical results providing the fast convergence guarantees for the iterate sequence generated by SAP along the top-$\ell$ dimensional subspace of $A$, with Theorem˜B.3 being the main result of this section. We use Theorem˜B.3 in the next section to derive the subspace convergence result for posterior mean in GP inference, proving Theorem˜3.2.